OpenStudy (zzelinski):

let f(x)=x^2-16 and g(x)=14-x Find (F-G)(9)

OpenStudy (anonymous):

F(9) = 81-16 G(9) = 14-9 (81-16) - (14-9)

OpenStudy (zzelinski):

Thanks!

OpenStudy (anonymous):

Anytime :) Medal? :D

OpenStudy (zzelinski):

yep